OpenStudy (anonymous):

One more question for today... Find the equation of the tangent line in Cartesian coordinates of the curve given polar coordinates by r=3sin(theta) at (theta)=pi/3

OpenStudy (anonymous):

oh wait i should read sorry

OpenStudy (anonymous):

that was totally wrong. i deleted it.

OpenStudy (anonymous):

can't you take the derivative, then convert to cartesian?

OpenStudy (anonymous):

graph is a circle with center at (1.5,0) and radius 1.5 yes?

OpenStudy (anonymous):

Lets take the pts where tangent intersect the circle (x,y) Now eq turns out be x^2 + y^2 >= (radius)^2 Here, radius = 3/2

OpenStudy (anonymous):

i'm with you so far but this circle is \[x^2+(y-\frac{3}{2})^2=\frac{9}{4}\] i htink

OpenStudy (anonymous):

center (0,3/2) radius 3/2

OpenStudy (anonymous):

bt how that y-3/2 comes up?

OpenStudy (anonymous):

\[\sqrt{x ^{2}+ y ^{2}} = R\]

OpenStudy (anonymous):

I did tht wrong sry.........3.sin pi/3 should be 3.(3)^1/2 ---- 2

OpenStudy (anonymous):

(radius)^2 should be 27/4

OpenStudy (anonymous):

this graph is in polar form yes? so when you graph \[r=\sin(\theta) \] it traces out a circle starting at (0,0) going up to (1,0) and back down to (0,0)

OpenStudy (anonymous):

bt its centre shoul be 0,0......am i right? satellite

OpenStudy (anonymous):

the graph is in polar form, not rectangular. the rectangular form of this circle, unless i really messed up somewhere, is \[x^2+(y-\frac{3}{2})^2=\frac{9}{4}\] a circle with radius [\frac{3}{2}\] and center at \[(0,\frac{3}{2})\] i am fairly certain of this but let me get a book

OpenStudy (anonymous):

a circle with center (0,0) and radius 1 would be \[r=1\]

OpenStudy (anonymous):

okay i might be wrong i haven't done question of polar with circle..... :)

OpenStudy (anonymous):

ok i have my book open. the idea is not to switch to rectangular. the idea is that \[\frac{dy}{dx}=\frac{\frac{dy}{d\theta}}{\frac{dx}{d\theta}}\]

OpenStudy (anonymous):

which book is tht ?

OpenStudy (anonymous):

which turns out to be \[\frac{\frac{dr}{d\theta}\sin(\theta)+r\cos(\theta)}{\frac{dr}{d\theta}\cos(x)-r\sin(\theta)}\]

OpenStudy (anonymous):

that is what we need to compute

OpenStudy (anonymous):

garden variety calc book. and i made a typo above. that \[\cos(x)\] should be \[\cos(\theta)\]

OpenStudy (anonymous):

now just plug in the numbers and get the answer

OpenStudy (anonymous):

but y do we differentiate everythng is constnt

OpenStudy (anonymous):

\[\frac{dr}{d\theta}=3\cos(\theta)\] and if \[\theta=\frac{\pi}{3}\] then \[r=\frac{\sqrt{3}}{2}\]

OpenStudy (anonymous):

damn another typo. this is annoying to write

OpenStudy (anonymous):

\[r=\frac{3\sqrt{3}}{2}\]

OpenStudy (anonymous):

so we get a numerator of \[3\cos(\theta)3\sin(\theta)+\frac{3\sqrt{3}}{2}\cos(\theta)\]

OpenStudy (anonymous):

and a denominator of \[3\cos^2(\theta) -\frac{3\sqrt{3}}{2}\sin(\theta)\]

OpenStudy (anonymous):

now put \[\theta =\frac{\pi}{3}\] to get the answer